LOG-IN
Displaying reports 1761-1780 of 3375.Go to page Start 85 86 87 88 89 90 91 92 93 End
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 13:55, Thursday 05 December 2019 (1912)Get code to link to this report
Seismic noise results from trillium seismomenter placed on TAMA floor
 
[Irene, Camilla, Federico]
 
Some conclusion drawn from data taken with a trillium seismomenter placed on TAMA floor (behind squeezing clean room, below mode cleaner pipe).
 
1) Figures 1, 2, 3 show spectrograms (GPS is in UTC). Only for the day 21/11 we have 24h of data.
 
2) Noise between 1-30 Hz changes between day and night. Figure 4 highlights it.
 
3) There is a device that turns off/off every ~ 8 minutes, which produces the peaks at 24.5 Hz (fan? air conditioning? pump?). See plot 5.
 
4) Figure 6 shows the percentiles calculated on a day. Some other non-stationary peaks, in addition to that at 24.5 Hz, are clearly visible
 
5) Comparing trillium spectrum with the TEAC seismometer we have in TAMA it seems the TEAC is useful only between 10Hz and 50 Hz. Its own noise dominates below 10Hz, while ADC noise dominates above 50Hz (To Be Confirmed) - Plot to be posted.
 
 
Good face to face!Figures 1, 2, 3 are spectrograms (GPS is in UTC). Only for the
November 21 we have 24h of data.
 
1) you see the peak around 10 Hz that you told us about. In reality
it is very wide, between 1-30 Hz, and from the spectrograms we see that it "breathes"
between day and night. Figure 4 highlights it, they are two spectra:
one by day, the other by night.
 
2) There is a device that turns off / on every 8 minutes or so, which makes the
peaks at 24.5 Hz (one fan? air conditioning? pump?). See plot 5.
 
3) Figure 6 is the percentiles calculated on a day. They look good
some non-stationary peaks, beyond that at 24.5 Hz.
 
That's all. We have no more precise conclusions, it should be understood
which are the devices that generate the peaks, but we have had no way
to analyze the data in time to do so. :)
 
Good face to face!
Images attached to this report
1912_20191205054737_191120trilliumzspectrogram.png 1912_20191205054743_191121trilliumzspectrogram.png 1912_20191205054749_191122trilliumzspectrogram.png 1912_20191205054912_seismometerdayvsnightfftz.png 1912_20191205054946_peak24hz.png 1912_20191205055022_percentiles21nov.png
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 13:22, Thursday 05 December 2019 (1911)Get code to link to this report
FC lock not stable

[Yuhang, Matteo, Eleonora]

On Tue 3/12 morning we found the FC has difficulties to acquire the lock and to keep it, while in the past months lock was very robust and easy to get.

After some investigation (and having tweeked the gain of ramepeato) we found the input power was reduced from 12 mW to 9 mW, due to temperature change of SHG.

The day before the air conditiong was switched from summer to winter mode by Yuhang.

The injected power was put back to 12mW and the gains were tuned to get UGF ~ 15 kHz.  (See pic 1 and 2). Anyway it seems a loop oscillation happened from time to time and error signal looks noisier than usual. It was quite late and we coudn't keep investingating any longer. We haven't been in TAMA yesterday and today because of F2F.  Will check the situation tomorrow.

Images attached to this report
1911_20191205051708_tf.jpeg 1911_20191205051734_gain.jpeg
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 11:47, Thursday 05 December 2019 (1910)Get code to link to this report
Comment to Mode mis-matching between filter cavity(FC) and local oscillator(LO) (Click here to view original report: 1905)

The main effect causing degradation of the overlap between LO and reflected beam form FC when it is ON and OFF resonance is the relative phase change between TEM 00 (resonating the cavity) and HOM (not resoanting). Raffaele did a quick calculation to show this and the numbers he found are in rough agreement with those we observed. I attached a pdf with a similar calulation and a plot that shows the overlap degradation as a function of the not coupled power into FC. 

Non-image files attached to this comment
R&D (General)
Print this report.
EleonoraCapocasa - 11:35, Thursday 05 December 2019 (1909)Get code to link to this report
FC dithering loops implemented for YAW

[Eleonora, Matteo, Raffaele]

On 29/11 dIthering was implemented also in Yaw, with the same scheme used for pitch.

Dithering lines are injected on INPUT and END mirrors at a frequency of 16.5 Hz and 17.5 Hz respectively. Each error signal obtained from transmitted power demodulation is filtered with a simple integrator and fed back to its own mirror.

From the time taken by the loop to recover the alignemnt after misalignig a mirror on purpose, the UGF seems smaller 50 mHz. I tried to measure openloop TF with a swep sine injection at very low frequency but coherence was not good, no matter the excitation amplitude.

Note that with the current filter configuration (simple integrator) the UGF of dithering is limited by the interaction with the damping loops. See attached simulation. We discussed the possibiliy to shape the filter to avod the instabilty. Anyway since damping UGF is quite high ~1 Hz. in the present configuration the sum of the two loops is stable. Another possibility would be to go back to DC controlled optical lever and add the dithering DC correction to the set point of this loop. I think this is what is done in Virgo.

Images attached to this report
1909_20191205033454_tfcross.png
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 18:56, Tuesday 03 December 2019 (1908)Get code to link to this report
Log on Dec. 3rd

I put fused silica mirrors inside the cryostat chamber in order to construct a FP cavity with cavity length od 9.9cm.
Then I injected a laser into the cavity and tried to see the flash with PD which is located at transmission port.
So far I could not see any transmitted beam.

I will improve the alignment and mode matching with two lenses.

R&D (FilterCavity)
Print this report.
YuhangZhao - 10:11, Tuesday 03 December 2019 (1907)Get code to link to this report
FDS (anti-squeezing~9dB rotates to squeezing~1dB in filter cavity) @ 2kHz

Yaochin and Yuhang

We see anti-squeezing rotates to squeezing @2kHz. But the filter cavity was really unstable.

average: 50 times

Setting of hand-fit parameters:

sqz_dB = 9.5;                      % produced SQZ

L_rt = 105e-6;                    % FC losses

L_inj = 0.17;                     % Injection losses

L_ro = 0.1;                      % Readout losses

A0 = 0.05;                         % Squeezed field/filter cavity mode mismatch losses

C0 = 0.03;                        % Squeezed field/local oscillator mode mismatch losses

ERR_L =   5e-12;                   % Lock accuracy [m]

ERR_csi = 50e-3;                  % Phase noise[rad]

phi_Hom = [-1/180*pi, -11/180*pi, -75/180*pi];         % Homodyne angle [rad] (you can input a vector of values)

det =  2e3;                       % detuning frequency 

int = 0.8e3;                         % frequency range = det+/-int 

 

Setting of shot noise level/detuning error parameters:

 

semilogx(fd+34, d_a1+133.5,'k','LineWidth',3) %squeezing to anti-squeezing

hold on

semilogx(fd+32, d_a2+133.5,'LineWidth',3) %anti-squeezing to squeezing

hold on

semilogx(fd+150, d_a3+133.5,'LineWidth',3) %intermediate

Images attached to this report
1907_20191203021515_fds.png
R&D (FilterCavity)
Print this report.
NaokiAritomi - 15:06, Saturday 30 November 2019 (1906)Get code to link to this report
FDS on 20191129

[Aritomi, Yuhang, Yaochin, Eleonora]

First we measured frequency independent squeezing with filter cavity (Pic. 1). We had 3.5dB squeezing down to 80Hz with filter cavity.

Then we alined BAB to AMC with 1kHz filter cavity detuning and measured FDS (Pic. 2). Fitting parameters are follows. It seems that detuning changed for squeezing and anti squeezing measurement.

CC2 demodulation phase
sqz: 75 deg, asqz: 130 deg, intermediate: 100 deg
 

sqz_dB = 13;                      % produced SQZ

L_rt = 100e-6;                    % FC losses

L_inj = 0.20;                     % Injection losses

L_ro = 0.11;                      % Readout losses

A0 = 0.1;                         % Squeezed field/filter cavity mode mismatch losses

C0 = 0.1;                         % Squeezed field/local oscillator mode mismatch losses

ERR_L =   5e-12;                  % Lock accuracy [m]

ERR_csi = 80e-3;                  % Phase noise[rad]

phi_Hom = [-5/180*pi, pi/2];      % Homodyne angle [rad] 

det =  [1074 1004];               % detuning frequency [Hz]

cavity_pole = 59.6;               % cavity pole (Hz)

 

We measured FDS again with smaller frequency span (Pic. 3). Fitting parameters are follows. This time, it seems that produced SQZ also changed in addition to detuning.

 

sqz_dB = [13.5 11 13.5];          % produced SQZ

L_rt = 100e-6;                    % FC losses

L_inj = 0.20;                     % Injection losses

L_ro = 0.11;                      % Readout losses

A0 = 0.05;                        % Squeezed field/filter cavity mode mismatch losses

C0 = 0.1;                         % Squeezed field/local oscillator mode mismatch losses

ERR_L =   5e-12;                  % Lock accuracy [m]

ERR_csi = 80e-3;                  % Phase noise[rad]

phi_Hom = [-3/180*pi 70/180*pi 7/180*pi];       % Homodyne angle [rad] 

det =  [970 1005 965];                          % detuning frequency [Hz]

cavity_pole = 59.6;               % cavity pole (Hz)

Images attached to this report
1906_20191130070637_sqz20191129.png 1906_20191130070643_fds20191129.png 1906_20191130074217_fds20191129expand.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 02:19, Saturday 30 November 2019 (1905)Get code to link to this report
Mode mis-matching between filter cavity(FC) and local oscillator(LO)

Aritomi, Eleonora, Yaochin, and Yuhang

For the adjustment of OPO/FC matching and FC/LO matching, we always do the following.

1. Send BAB into the filter cavity. Align IR into the filter cavity while filter cavity alignment is kept in the best condition. Check IR 0/1/2 order modes by changing AOM frequency.

2. Align BAB reflection into AMC while the filter cavity is locked.

However, we found that the matching from BAB_ref into AMC is quite different when the filter cavity is locked and detuned. And we know that we measure FDS while FC is detuned, so we decide to align BAB_ref into AMC while FC is detuned. So now the procedure is as follows.

1. Send BAB into the filter cavity. Align IR into the filter cavity while filter cavity alignment is kept in the best condition. Check IR 0/1/2 order modes by changing AOM frequency.

2. Align BAB reflection into AMC while the filter cavity is detuned.

However, we are not very sure why the matching can be so different when FC is locked and detuned. Because I think the BAB_ref should be quite similar in both cases of lock and detuned since we had 94% of matching from OPO/FC. I put the reason for my thought as following. Please correct me if there is something wrong.

Images attached to this report
1905_20191129181900_wechatimg602.jpeg
Comments related to this report
EleonoraCapocasa - 11:47, Thursday 05 December 2019 (1910)

The main effect causing degradation of the overlap between LO and reflected beam form FC when it is ON and OFF resonance is the relative phase change between TEM 00 (resonating the cavity) and HOM (not resoanting). Raffaele did a quick calculation to show this and the numbers he found are in rough agreement with those we observed. I attached a pdf with a similar calulation and a plot that shows the overlap degradation as a function of the not coupled power into FC. 

Non-image files attached to this comment
R&D (FilterCavity)
Print this report.
YaoChinHuang - 01:04, Saturday 30 November 2019 (1904)Get code to link to this report
IR light jitter on LO & FC path

Yao-Chin and Yuhang

 

When IRMC locking, we monitored yaw & pitch of LO light jittering by position sensitive detector(PSD) as shown in fig 1. PSD was put before alignment mode cavity(AMC). We also measured yaw & pitch of PSD dark noise.

In addition, we sent 200mV Vpp of random noise to IR phase shifter(IRPS) and measured again yaw & pitch of LO light jittering as shown in fig 2. Magnitude value increased more obviously above 60Hz than without random noise situation.

When BAB light sent to filter cavity, we monitored yaw & pitch of reflect light by PSD as shown in fig 3. Because suspension mirror jittering, The jittering magnitude of low frequency range (<200Hz for yaw, <500Hz for pitch) is huge.

Images attached to this report
1904_20191129163406_figure1.png 1904_20191129163426_figure2.png 1904_20191129163437_figure3.png
KAGRA MIR (Absorption)
Print this report.
PengboLi - 10:29, Friday 29 November 2019 (1902)Get code to link to this report
Result of the beam profile measured by the beam profiler

Simon, Pengbo

We finished the measurement of the beam profile using the beam profiler.

The beam profile could not be measured at the waist due to the space limits, and the results may be not that accurate.

But as can be seen from the attachment, the diameter of the waist is 81 mum roughly, which didn't offset very much compared with the result before. 

Images attached to this report
1902_20191129022346_figure.png
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 17:39, Thursday 28 November 2019 (1901)Get code to link to this report
Glueing Jig

I designed a glueing jig for fused silica mirrors.
There is a through hole for putting a mirror and its holder.
A space aroung the center is in order to prevent sticking the jig to the mirror due to the leaking glue.

I will ask Sato-san to check the design.

Non-image files attached to this report
R&D (FilterCavity)
Print this report.
NaokiAritomi - 17:00, Thursday 28 November 2019 (1900)Get code to link to this report
FDS after DC balance on 20191126

sqz_dB = 12.5;                    % produced SQZ

L_rt = 100e-6;                    % FC losses

L_inj = 0.20;                     % Injection losses

L_ro = 0.11;                      % Readout losses

A0 = 0.1;                         % Squeezed field/filter cavity mode mismatch losses

C0 = 0.1;                         % Squeezed field/local oscillator mode mismatch losses

ERR_L =   5e-12;                  % Lock accuracy [m]

ERR_csi = 80e-3;                  % Phase noise[rad]

phi_Hom = [0, pi/2*105/90];       % Homodyne angle [rad] (you can input a vector of values)

det =  360;                       % detuning frequency 

Images attached to this report
1900_20191129114114_20191126rot.png
R&D (FilterCavity)
Print this report.
YaoChinHuang - 01:56, Thursday 28 November 2019 (1899)Get code to link to this report
IR phase shifter with random noise respond on homodyne detector

Yao-Chin and Yuhang

 

When IRMC locking and only LO light hit into homodyne detector, we aligned LO light to arrive dc balance signal of homodyne shown in fig 1 and measured shot noise using analyzer (Aglient 35670A). Using low frequency, which range from 125mHz to 200Hz, random noise from analyzer source sent to high voltage amplifier (Gain: x15) of IR phase shifter(CC2). With different Vpp values of random noise (its spectrum shown in fig 2), we measured the RF spectrums of homodyne detector. Below 100mV Vpp, the spectrums are similar from 20Hz to 200Hz range shown in fig 3. The magnitude value of shot noise increased when above 500mV Vpp of random noise. 

Vpp of random noise

(mV)

Average value

(dB)

0 -133.40
5 -133.34
10 -133.40
50 -133.34
100 -133.29
500 -128.40
1000 -120.98

Note: We average spectrum magnitude value from 20Hz to 200Hz.

Images attached to this report
1899_20191127172112_img3351.jpg 1899_20191127172131_figure2.png 1899_20191127174416_figure1.png
KAGRA MIR (Absorption)
Print this report.
SimonZeidler - 00:30, Thursday 28 November 2019 (1898)Get code to link to this report
Results of beam-profile measurements around the sample holder

Pengbo, Simon

We finished the measurements of the beam-profile after the telescope.
The horizontal profile could not be measured directly at the waist (as we put the blade too much on the sensor-side). However, the fit with a Gaussian diffraction development gives a good result in terms of the expected waist diameter of 70 mum (~ 72 mum from the fit, see attached picture).

For the vertical measurements, we put the blade ~2cm closer to the pump-beam side and started another run. This time (see the second figure attached) we could cross the beam's waist in the expected position which is in correlation with the horizontal measurements. Also here, the distribution has been fitted and the diameter estimated to be ~74 mum at the waist, also in agreement with the horizontal measurements.

Images attached to this report
1898_20191127163031_figure1.png 1898_20191128024553_figure11.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 21:58, Wednesday 27 November 2019 (1895)Get code to link to this report
IR locking accuracy after the improvement of BAB matching into filter cavity

Eleonora, Raffaele, Yaochin and Yuhang

Yesterday night we came to TAMA and improved especially the matching and alignment from OPOtra to filter cavity. We moved the lens position and also tried to align a bit the steering mirror. After this work, the matching level is estimated as following

TEM00: 460

TEM10: 108

LG10: 108

offset: 100

So the mis-matching level is about 5%

This matching level is better than the measurement of locking accuracy of last time. And we think the matching level will also influence the locking accuracy. So we characterize the locking accuracy again. This time we used the same setting as last time. The AOM was scanned with a modulated sine wave. The modulation is 2000Hz/s. We measured the PDH signal and use it for calibration. The measured PDH signal pk-pk is 234mV with a seperation of 118ms.

So the calibration is 2000/(2*234/118) Hz/V

We measured the demodulated BAB reflection spectrum, the locking accuracy is integrated as 5Hz. But now we have some new peaks.

Images attached to this report
1895_20191127140031_figure1.png 1895_20191127140412_tek00080.png
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 20:30, Wednesday 27 November 2019 (1894)Get code to link to this report
Brief inspection of the spacer

Today, the spacer for cryogenic cavity was delivered and I did brief inspection of it with half inch flat mirror.
I just checked whether the reflected beam can pass through the output hole of the spacer or not.
The reflected beam could pass through, but there is a undesirable feature due to my mistake of design.
Fortunately it can be easily removed by an additional machining.
I will ask the company to do it.

Images attached to this report
1894_20191127123047_20191127spacer.jpg
R&D (FilterCavity)
Print this report.
YuhangZhao - 16:06, Wednesday 27 November 2019 (1893)Get code to link to this report
Characterization of CC2 loop after the replacement of new phase shifter

The measurement of CC2 loop opto-mechanical transfer function, CC2 open loop tranfer function and CC2 error signal spectrum is attached as attachement.

The good news is that the OMTF is very flat.

The first obvious oscillation we could find is ~23kHz(the last attachement). While we could see from the error singal spectrum, we have two narrow peaks at ~15.5kHz and ~22kHz.

Images attached to this report
1893_20191127080503_cc2omtf.png 1893_20191127080513_cc2oltf.png 1893_20191127080607_cc2err.png 1893_20191127081109_resonance.jpg
R&D (FilterCavity)
Print this report.
YuhangZhao - 15:39, Wednesday 27 November 2019 (1892)Get code to link to this report
green power change

Aritomi, Eleonora, Yaochin and Yuhang

We found that fc tra is more than 3000 counts. And this is more than we had, so we did the following check. And it shows that the more green power seems to be from a higher conversion efficiency of SHG.

green before EOM 268mW

before AOM 48.4mW

before MZ 198mW

AOM frequency 109.03575MHz

AOM amplitude 2.5dbm (then go through zhl-2 amplifier to AOM)

R&D (FilterCavity)
Print this report.
NaokiAritomi - 23:59, Tuesday 26 November 2019 (1891)Get code to link to this report
FDS at 600Hz with new phase shifter and testmass feedback

[Aritomi, Yuhang, Yaochin, Eleonora, Matteo, Raffaele]

First we aligned IR into filter cavity. Current mode matching is around 95.8%.

Mode IR transmission
TEM00 400
HG10 112
HG01 105
offset 102

Then we succeeded in locking CC2 with filter cavity with new phase shifter. This time CC2 testmass feedback worked well and CC2 correction signal and IRMC reflection became more stable. Gain of testmass feedback is -2.

We measured FDS at 600Hz and CC2 phase noise (Pic.1,2). CC2 demodulation phase is as follows. CC2 error signal is 76mVpp. At high frequency, we had 3dB squeezing, but the spectrum was not clean. Below 100Hz, there was large bump.

  CC2 demodulation phase for SQZ (deg) CC2 demodulation phase for ASQZ (deg)
FDS 70 100
FIS 95  

After this measurement, we found that DC balance of LO was bad. We measured frequency independent squeezing before/after DC balance (Pic.3). After DC balance, squeezing spectrum and squeezing level got better, but still not clean as before.

Then we measured FDS again with/without testmass feedback (Pic.4). Low frequency bump became lower down to 60Hz and it doesn't change with/without testmass feedback. It seems that DC balance improved low frequency bump.

Images attached to this report
1891_20191126155846_fds20191126.png 1891_20191126155854_cc220191126.png 1891_20191126155903_sqz20191126.png 1891_20191126155908_fdstm20191126.png
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 16:02, Tuesday 26 November 2019 (1890)Get code to link to this report
Transmittance of fused silica mirror

I roughly investigated the transmittance of fused silica mirrors which will be used for input and output couplers.
I measured transmitted beam power for 2 of 4 mirrors.
The trasmitted power was 1.8uW for both of them with respect to 10mW incoming power.
This value corresponds to T=0.018% and R=99.982% assuming no loss in mirrors.
Then the finesse can be estimated as 1.7*104.
This value is reasonable since the designed finesse was 1.5*104.

I gonna construct FP cavity with these mirrors and try to lock TEM00.